OpenStudy (anonymous):

3) At the low point in it swing, a pendulum bob with a mass of 0.19 kg has a velocity of 4.53 m/s. (a.) What is its kinetic energy at the low point? (b.) Ignoring air resistance, how high will the bob swing above the low point before reversing direction?

OpenStudy (anonymous):

Kinetic energy is given by, E= (1/2) m v^2 E= (1/2) (0.19)(4.53^2) = 1.949 = 1.95J approx. Loss in kinetic energy = Gain in potential energy 1.949 = mgh 1.949 = (0.19)(9.81)(h) h= 1.0459m = 1.05m approx.